A Spherical Tank with a radius .25m is filled with water (p=1000kg/m^3). Calculate the mass and weight of the water in SI units.: A Spherical Tank with a radius .25m is filled with water (p=1000kg/m^3). Calculate the mass and weight of the water in SI units.
Answer by orca(336)

The volume of a sphere with radius r can be found by the formula:
V=(4/3)*pi*r^3
So the volume V of the tank is:
V=(4/3)*pi*0.25^3=0.0654m^3
The mass of the water can be found by multiplying V by p.
0.0654m^3*1000(kg/m^3)=65.4kg